Write the standard form of the equation of the line that passes through the given points. Can you show me how you got this?
(5,8), (3,2)

Answer by mananth(16946)

Write the standard form of the equation of the line that passes through the given points. Can you show me how you got this?
(5,8), (3,2)
...
slope between these points = m= (y2-1)/(x2-x1)
m=(2-8)/3-5)
m=-6/-2
m=3
..
find the y intercept by plugging values of any of the points and the slope m
(3,2) m=3
2=3*3+b
2=9+b
b=-7
...
the equation will be y= 3x-7
y-3x=-7